Current Status

Current Status

3 September 2002: After some trouble this year with keeping UFTI cold, the instrument was warmed up, the O-ring in the entrance window replaced, and a leak in the cryostat patched. The instrument is now cold again and holding its temperature/vacuum well. – CJD

1 March 2002: After some trouble with the array controller, the “5e-/DN” unit was installed. i.e the gain has thus changed from 7.4e-/DN and now stands at 5.5e-/DN (details are given on the main parameters web page). – CJD

15 November 2001: Following some temperature-related problems, UFTI was warmed up in mid-November. The instrument is back on the telescope now and operating properly; the array orientation and pixel scales were updated on the web in early February 2002 (though these new values apply to all data taken after 15 November). – CJD

5 September 2001:. New bad-pixel mask implemented at UKIRT and the JAC. – CJD

11 April 2001: UFTI was today warmed up to replace the faulty 2-1S(1) filter and to check the condition of filter wheel one, which has been failing recently. The CO 1% (2.32um) filter was also replaced with a Shack Hartmann mask which we hope to use for focusing in the near future. See the UFTI Parameters web page for a list of available filters. – CJD

2 February 2001:. A new bad-pixel mask has been implemented at UKIRT and the JAC. – CJD

11 October 2000:. Due to the delayed delivery of a number of filters, we have unfortunately had to postpone the filter change that was scheduled for October 2000. – CJD

17 August 2000:. A new bad-pixel mask has been implemented at UKIRT and the JAC. – CJD

13 May 2000:. UFTI is on the telescope and working well. New filters were installed in spring 2000; all are ok except the 2-1S(1) filter, which unfortunately produces elongated images. This filter will be replaced later in the year with a “Mauna Kea Consortium” version (as will a few of the others). – CJD